Luther Corporation Consolidated Balance Sheet December 31, 2006 and 2005 (in $ millions) Assets 2006 2005 Liabilities and Stockholders' Equity 2006 2005 Current Assets ​ ​ Current Liabilities ​ ​ Cash 56.1 58.5 Accounts payable 88.1 73.5 Accounts receivable 54.5 39.6 Notes payable / short-term debt 10.9 9.6 Inventories 44.8​ 42.9 Current maturities of long-term debt 40.7 36.9 Other current assets 5.0 3.0 Other current liabilities 6.0 12.0 Total current assets 160.4​ 144.0 Total current liabilities 145.7​ 132.0 ​ ​ ​ ​ ​ ​ Long-Term Assets ​ ​ Long-Term Liabilities ​ ​ Land 66.8 62.1 Long-term debt 227​ 168.9 Buildings 108.5​ 91.5 Capital lease obligations ​ ​ Equipment 117.1​ 99.6 ​ ​ ​ Less accumulated depreciation (54.4) (52.5) Deferred taxes 22.8 22.2 Net property, plant, and equipment 238​ 200.7 Other long-term liabilities --- --- Goodwill 60.0 -- Total long-term liabilities 249.8​ 191.1 Other long-term assets 63.0 42.0 Total liabilities 395.5​ 323.1 Total long-term assets 361​ 242.7 Stockholders' Equity 125.9​ 63.6 ​ ​ ​ ​ ​ ​ Total Assets 521.4​ 386.7 Total liabilities and Stockholders' Equity 521.4​ 386.7 Refer to the balance sheet above. If in 2006 Luther has 10.2 million shares outstanding and these shares are trading at $16 per share, then what is Luther's enterprise value? Select one: A. -$540.0 million B. $771.4 million C. $385.7 million D. $521.4 million
Answer #1

The correct answer is option C. $385.7 million.

Enterprise Value = Market Value of Equity + Market Value of all types of debt - Cash on the balance sheet

Market Value of equity = Current share price x Nos. of shares outstanding = $ 16 x 10.20 mn = $ 163.20 mn

Market value of all types of debt = Market value of long term debt + market value of current portion of long term debt + Short term borrowing / notes payable.

Assumption: Market value of debts = Book value of debts

Hence, market value of debt = 227 + 40.7 + 10.9 = $ 278.60 mn

Cash on the balance sheet = $ 56.10 mn

Hence, Enterprise Value = Market Value of Equity + Market Value of all types of debt - Cash on the balance sheet = 163.20 + 278.60 - 56.10 = $ 385.70 mn

Hence, the correct answer is option C. $385.7 million.